OpenStudy (anonymous):

-2x+4=-2(3x+16)How do I solve this?

OpenStudy (tkhunny):

I would first divide by -2 - everything.

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you, I'm going to study this

OpenStudy (anonymous):

do you mean multiply?

OpenStudy (anonymous):

@jberlinghoff does it make sense? you want me to explain it?

OpenStudy (anonymous):

please explain. thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

o.k- Thank you! I will practice these! I'm sure I will be back.

OpenStudy (tkhunny):

-2x+4=-2(3x+16) Seriously, divide by -2, first. x - 2 = 3x + 16 It's about 10 times easier.